# Partial Rewriting and Value Interpretation of Logically Constrained Terms

### v1

@@ 1 Introduction
Many topics from the term rewriting literature have already been studied within the LCTRSs formalism, including such as confluence analysis [14, 16], (non-)termination analysis [7, 11], completion [19], rewriting induction [9, 4], algebraic semantics [1], and complexity analysis [20]. Nevertheless, the framework also exhibits aspects that are not covered by traditional term rewriting. One of such aspect is the rewriting of constrained terms. A constrained term consists of a term together with a constraint that restricts the ways in which the term may be instantiated. For example, the constrained term $\mathsf{f}(x)\penalty 10000\ [\hskip 0.85358ptx>2\hskip 0.85358pt]$ (in LCTRS notation) can be intuitively viewed as a set of terms $\{\hskip 0.85358pt\mathsf{f}(x)\mid x>2\hskip 0.85358pt\}$. In the analysis of LCTRSs, constrained rewrite rules must be handled, and consequently rewriting constrained terms naturally arises. Indeed, constrained rewriting is an integral part of many different analysis techniques. For instance, in confluence analysis, finding a specific joining sequence for constrained critical pairs often requires reasoning about two terms under a shared constraint, which leads to working on constrained terms. Similarly, in rewriting induction, rewriting of constrained terms is employed in several of its inference steps.
@@
Rewriting constrained terms was already introduced in early work on LCTRSs [8]. However, the original formulation of rewriting constrained terms is difficult to deal with, since rewrite steps are used modulo equivalence transformations of constrained terms, and the results of rewriting a constrained term remain highly flexible even when both the rewrite rule and the position are fixed. These difficulties motivated us to revisit the formalism of constrained rewriting in LCTRSs in our recent paper [takahataPPDP]. In that work, we have introduced the notions of existentially constrained terms and most general constrained rewriting for left-linear LCTRSs. We have shown the uniqueness of reducts in this new setting, as well as commutativity between rewrite steps and equivalence, under the restriction to left-value-free rewrite rules [^1] However, any rewrite rule is simulated by a left-value-free rewrite rule, and thus this is not a real restriction.. Our formalism of most general constrained rewriting extracts the so-called “most general” part of the corresponding original formalism. As a result, it offers several advantages over the original approach of rewriting constrained terms.
@@
In this paper, we introduce another notion of rewriting on existentially constrained terms, called partial constrained rewriting. The difference between the two notions can be explained informally as follows. On the one hand, most general constrained rewriting applies a rule to all instantiations of a constrained terms. On the other hand, partial constrained rewriting applies a rule to only some instantiations of a constrained term (see LABEL:fig:intuition_for_twokinds_of_rewriting).
@@
$\begin{array}[]{lclc}\mathsf{sum}(x)\penalty 10000\ [1\leqslant x\land x\leqslant 5]&\to_{\mathcal{R}}&x+\mathsf{sum}(x+-1)\penalty 10000\ [1\leqslant x\land x\leqslant 5]\\ &\to_{\mathcal{R}}&x+\mathsf{sum}(y)\penalty 10000\ [1\leqslant x\land x\leqslant 5\land y=x-1]$$\not\to_{\mathcal{R}}\end{array}$
obtains a normal form of most general constrained rewriting. On the other hand, we have $\mathsf{sum}(1)\to_{\mathcal{R}}^{*}1$, $\mathsf{sum}(2)\to_{\mathcal{R}}^{*}3$, $\mathsf{sum}(3)\to_{\mathcal{R}}^{*}6$, $\mathsf{sum}(4)\to_{\mathcal{R}}^{*}10$, $\mathsf{sum}(5)\to_{\mathcal{R}}^{*}15$. However, most general rewriting is not suitable for this computation, as there is no single rule that can be applied to the constrained term $\mathsf{sum}(y)\penalty 10000\ [1\leqslant x\land x\leqslant 5\land y=x-1]$. In contrast, partial constrained rewriting allows the rewriting process to proceed by
@@
$\begin{array}[]{lcl}\cdots&\to_{\mathcal{R}}&1+\mathsf{sum}(y)\penalty 10000\ [1\leqslant x\land x\leqslant 5\land y=x-1]\\ &\leadsto_{\mathcal{R}}&1+(1+\mathsf{sum}(z))\penalty 10000\ [1\leqslant x\land x\leqslant 5\land y=x-1\land z=y-1]\end{array}$
Thus, by partial rewriting we obtain $\{t\in\mathsf{NF}\mid\mathsf{sum}(x)\penalty 10000\ [1\leqslant x\land x\leqslant 5]\leadsto^{*}t\}=\{1,3,6,10,15\}$.
@@
In this paper, we study the differences between most general rewriting and partial rewriting from several perspectives. First, we establish a direct correspondence between the two formalisms, showing that partial constrained rewrite steps are included in most general rewrite steps, but not vice versa. We furhter demonstrate that most general rewrite steps can be simulated by partial rewriting through subsumption and equivalence of constrained terms. We also provide characterizations of each type of rewriting constrained terms in terms of interpretations of constrained terms and rewrite rules as sets of terms and rewrite rules, respectively. The idea of interpreting (original) constrained terms as sets of terms has appeared, for example, in [4, Section 2.3] to reduce certain problems on LCTRSs to problems in TRSs. However, these characterizations alone are insufficient to distinguish partial constrained reductions from most general constrained reductions. This motivates the introduction of a novel notion of value interpretation. Like the standard interpretation, the value interpretation reflects the equivalence of constrained terms; however, unlike the standard interpretation, it does not reflect subsumption. Using the value interpretation, we provide more precise characterizations of the two notions of rewriting constrained terms, highlighting subtle differences between partial and most general rewriting.
@@
The remainder of the paper is organized as follows. Section 2 introduces some preliminaries. In Section 3, we introduce partial constrained rewriting and establish direct correspondences between partial constrained rewriting and most general constrained rewriting. Section 4 explores properties of interpretations of existentially constrained terms, introduces the interpretation of constrained rewrite rules, and characterizes both partial and most general rewrite steps based on these interpretations, while also highlighting the limitations of these. In Section 5, we introduce the value interpretation and discuss its properties. Section 6 provides characterizations of partial and most general rewrite steps using value interpretations. Related work is discussed in Section 7, and Section 8 concludes the paper.
@@ 2 Preliminaries
In this section, we briefly recall the basic notions of LCTRSs [1, 8, 14, 16, 17] and fix additional notations used throughout this paper. Familiarity with the basic notions of term rewriting is assumed (see, e.g., [2]).
@@
We employ a sorted signature $\langle\mathcal{S},\mathcal{F}\rangle$ consisting of a set $\mathcal{S}$ of sorts and a set $\mathcal{F}$ of function symbols, where each $f\in\mathcal{F}$ is attached with its sort declaration $f\colon{\tau_{1}}\times\cdots\times{\tau_{n}}\to\tau_{0}$ ($\tau_{0},\ldots,\tau_{n}\in\mathcal{S}$). The set of sorted variables is denoted by $\mathcal{V}$. For $X\subseteq\mathcal{V}$ and $F\subseteq\mathcal{F}$, the set of term over the function symbols in $F$ and variables in $X$ is denoted by $\mathcal{T}(F,X)$. We write $t^{\tau}$ to denote that a term $t$ has sort $\tau$. The set of variables appearing in a term $t$ is denoted by $\mathcal{V}\mathsf{ar}(t)$. A term $t$ is said to be linear when any variable occurs at most once in $t$. For brevity, we write a sequence of variables as $\vec{x}$, etc. The set of variables occurring in $\vec{x}$ is denoted by $\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}$. Likewise, the set of sequences of elements of a set $T$ is denoted by $T^{*}$.
@@
The set of positions in a term $t$ is denoted by $\mathcal{P}\mathsf{os}(t)$. The symbol and subterm occurring at a position $p\in\mathcal{P}\mathsf{os}(t)$ is denoted by $t(p)$ and $t|_{p}$, respectively. A term obtained from $t$ by replacing subterms at positions ${p_{1}},\dots,{p_{n}}$ by the terms ${t_{1}},\dots,{t_{n}}$, having the same sort as $t|_{p_{1}},\ldots,t|_{p_{n}}$, as $t[{t_{1}},\dots,{t_{n}}]_{{p_{1}},\dots,{p_{n}}}$ or just $t[{t_{1}},\dots,{t_{n}}]$ when no confusions arises. Sometimes we consider an expression obtained by replacing those subterms $t|_{p_{1}},\ldots,t|_{p_{n}}$ in $t$ by holes of the same sorts, which is called a multihole context and denoted by $t[\penalty 10000\ ]_{{p_{1}},\dots,{p_{n}}}$. A sort-preserving function $\sigma$ from $\mathcal{V}$ to $\mathcal{T}(\mathcal{F},\mathcal{V})$ whose domain $\mathcal{D}\mathsf{om}(\sigma)=\{\hskip 0.85358ptx\in\mathcal{V}\mid\sigma(x)\neq x\hskip 0.85358pt\}$ is finite is called a substitution, where it is identified with its homomorphic extension $\sigma\colon\mathcal{T}(\mathcal{F},\mathcal{V})\to\mathcal{T}(\mathcal{F},\mathcal{V})$. For a set of terms $U$, we write $\sigma(U)=\{\hskip 0.85358pt\sigma(t)\mid t\in U\hskip 0.85358pt\}$. As usual, $\sigma(t)$ is also written as $t\sigma$. We write $\{x_{1}:=t_{1},\ldots,x_{n}:=t_{n}\}$ for the substitution $\sigma$ such that $\sigma(x_{i})=t_{i}$ ($1\leqslant i\leqslant n)$ and $\mathcal{D}\mathsf{om}(\sigma)\subseteq\{x_{1},\ldots,x_{n}\}$, which also be written as $\{\vec{x}:=\vec{t}\}$ for brevity.
@@
We assume that the set $\mathcal{S}$ of sorts and the set $\mathcal{F}$ of function symbols are partitioned into two disjoint sets, respectively: $\mathcal{S}=\mathcal{S}_{\mathsf{th}}\uplus\mathcal{S}_{\mathsf{te}}$ and $\mathcal{F}=\mathcal{F}_{\mathsf{th}}\uplus\mathcal{F}_{\mathsf{te}}$, where each $f\colon{\tau_{1}}\times\cdots\times{\tau_{n}}\to\tau_{0}\in\mathcal{F}_{\mathsf{th}}$ satisfies $\tau_{i}\in\mathcal{S}_{\mathsf{th}}$ for all $0\leqslant i\leqslant n$. Elements of $\mathcal{S}_{\mathsf{th}}$ ($\mathcal{F}_{\mathsf{th}}$) and $\mathcal{S}_{\mathsf{te}}$ ($\mathcal{F}_{\mathsf{te}}$) are called theory sorts (symbols) and term sorts (symbols). Moreover, we consider a model $\mathcal{M}=\langle\mathcal{I},\mathcal{J}\rangle$ over the sorted signature $\langle\mathcal{S}_{\mathsf{th}},\mathcal{F}_{\mathsf{th}}\rangle$. Here, $\mathcal{I}$ is interpretations for theory sorts, and $\mathcal{I}$ assigns any sort $\tau\in\mathcal{S}_{\mathsf{th}}$ a non-empty set $\mathcal{I}(\tau)$; $\mathcal{J}$ is interpretations for theory symbols, and $\mathcal{J}$ assigns any theory symbol $f\colon{\tau_{1}}\times\cdots\times{\tau_{n}}\to\tau_{0}\in\mathcal{F}_{\mathsf{th}}$ to a function $\mathcal{J}(f)\colon\mathcal{I}(\tau_{1})\times\cdots\times\mathcal{I}(\tau_{n})\to\mathcal{I}(\tau_{0})$. A valuation $\rho$ on the model $\mathcal{M}=\langle\mathcal{I},\mathcal{J}\rangle$ is a mapping that assigns any $x^{\tau}\in\mathcal{V}$ with $\rho(x)\in\mathcal{I}(\tau)$. The interpretation of a term $t$ in the model $\mathcal{M}$ over the valuation $\rho$ is denoted by ${t}]\!]_{\mathcal{M},\rho}$.
@@
We further assume that any element $a\in\mathcal{I}(\tau)$ appears as a constant $a^{\tau}$ in $\mathcal{F}_{\mathsf{th}}$, for every $\tau\in\mathcal{S}_{\mathsf{th}}$. These constants are called values and the set of all values is denoted by $\mathcal{V}\mathsf{al}$. The set of values contained in a term $t$ is denoted by $\mathcal{V}\mathsf{al}(t)$. A term $t$ is value-free when $\mathcal{V}\mathsf{al}(t)=\varnothing$. For a substitution $\gamma$, we define $\mathcal{V}\mathcal{D}\mathsf{om}(\gamma)=\{\hskip 0.85358ptx\in\mathcal{V}\mid\gamma(x)\in\mathcal{V}\mathsf{al}\hskip 0.85358pt\}$.

### v2

@@ 1 Introduction
Many topics from the term rewriting literature have already been studied within the LCTRS formalism, including such as confluence analysis [14, 16], (non-)termination analysis [7, 11], completion [19], rewriting induction [9, 4], algebraic semantics [1], and complexity analysis [20]. Nevertheless, the framework also exhibits aspects that are not covered by traditional term rewriting. One of such aspect is the rewriting of constrained terms. A constrained term consists of a term together with a constraint that restricts the ways in which the term may be instantiated. For example, the constrained term $\mathsf{f}(x)\penalty 10000\ [\hskip 0.85358ptx>2\hskip 0.85358pt]$ (in LCTRS notation) can be intuitively viewed as a set of terms $\{\hskip 0.85358pt\mathsf{f}(x)\mid x>2\hskip 0.85358pt\}$. In the analysis of LCTRSs, constrained rewrite rules must be handled, and consequently rewriting constrained terms naturally arises. Indeed, constrained rewriting is an integral part of many different analysis techniques. For instance, in confluence analysis, finding a specific joining sequence for constrained critical pairs often requires reasoning about two terms under a shared constraint, which leads to working on constrained terms. Similarly, in rewriting induction, rewriting of constrained terms is employed in several of its inference steps.
@@
Rewriting constrained terms was already introduced in early work on LCTRSs [8]. However, the original formulation of rewriting constrained terms is difficult to deal with, since rewrite steps are used modulo equivalence transformations of constrained terms, and the results of rewriting a constrained term remain highly flexible even when both the rewrite rule and the position are fixed. These difficulties motivated us to revisit the formalism of constrained rewriting for LCTRSs in [18]. In that work, we have introduced the notions of existentially constrained terms and most general constrained rewriting for left-linear LCTRSs. We have shown the uniqueness of reducts in this new setting, as well as commutativity between rewrite steps and equivalence, under the restriction to left-value-free rewrite rules [^1] However, any rewrite rule is simulated by a left-value-free rewrite rule, and thus this is not a real restriction.. Our formalism of most general constrained rewriting extracts the so-called “most general” part of the corresponding original formalism. As a result, it offers several advantages over the original approach of rewriting constrained terms.
@@
In this paper, we introduce another notion of rewriting on existentially constrained terms, called partial constrained rewriting. The difference between the two notions can be explained informally as follows. On the one hand, most general constrained rewriting applies a rule to all instantiations of a constrained term. On the other hand, partial constrained rewriting applies a rule to only some instantiations of a constrained term (see LABEL:fig:intuition_for_twokinds_of_rewriting).
@@
Let us consider the following LCTRS over the integer theory:
@@
The following reduction obtains a normal form by most general constrained rewriting:
[…]
On the other hand, we have $\mathsf{sum}(1)\to_{\mathcal{R}}^{*}1$, $\mathsf{sum}(2)\to_{\mathcal{R}}^{*}3$, $\mathsf{sum}(3)\to_{\mathcal{R}}^{*}6$, $\mathsf{sum}(4)\to_{\mathcal{R}}^{*}10$, and $\mathsf{sum}(5)\to_{\mathcal{R}}^{*}15$. However, most general rewriting is not suitable for this computation, as there is no single rule that can be applied to the constrained term $\mathsf{sum}(y)\penalty 10000\ [1\leqslant x\land x\leqslant 5\land y=x-1]$. In contrast, partial constrained rewriting allows the rewriting process to proceed by
@@
[…]
Thus, by partial rewriting we obtain $\{\hskip 0.85358ptt\in\mathsf{NF}\mid\mathsf{sum}(x)\penalty 10000\ [1\leqslant x\land x\leqslant 5]\leadsto^{*}t\hskip 0.85358pt\}=\{\hskip 0.85358pt1,3,6,10,15\hskip 0.85358pt\}$.
@@
In this paper, we study the differences between most general rewriting and partial rewriting from several perspectives. First, we establish a direct correspondence between the two formalisms, showing that partial constrained rewrite steps are included in most general rewrite steps, but not vice versa. We further demonstrate that most general rewrite steps can be simulated by partial rewriting through subsumption and equivalence of constrained terms. We also provide characterizations of each type of rewriting constrained terms in terms of interpretations of constrained terms and rewrite rules as sets of terms and rewrite rules, respectively. The idea of interpreting constrained terms as sets of terms has appeared, for example, in [4, Section 2.3] to reduce certain problems on LCTRSs to problems in TRSs. However, these characterizations alone are insufficient to distinguish partial constrained reductions from most general constrained reductions. This motivates the introduction of a novel notion of value interpretation. Like the standard interpretation, the value interpretation reflects the equivalence of constrained terms; however, unlike the standard interpretation, it does not reflect subsumption. Using the value interpretation, we provide more precise characterizations of the two notions of rewriting constrained terms, highlighting subtle differences between partial and most general rewriting.
@@
The remainder of the paper is organized as follows. Section˜2 introduces some preliminaries. In Section˜3, we introduce partial constrained rewriting and establish direct correspondences between partial constrained rewriting and most general constrained rewriting. Section˜4 explores properties of interpretations of existentially constrained terms, introduces the interpretation of constrained rewrite rules, and characterizes both partial and most general rewrite steps based on these interpretations, while also highlighting the limitations of these. In Section˜5, we introduce the value interpretation and discuss its properties. Section˜6 provides characterizations of partial and most general rewrite steps using value interpretations. Related work is discussed in Section˜7, and Section˜8 concludes the paper. All omitted proofs of the claims are provided in the appendix.
@@ 2 Preliminaries
In this section, we briefly recall the basic notions of LCTRSs [1, 8, 14, 16, 17] and fix additional notations used throughout this paper. Familiarity with the basic notions of term rewriting is assumed (see for example [3]).
@@
We employ a sorted signature $\langle\mathcal{S},\mathcal{F}\rangle$ consisting of a set $\mathcal{S}$ of sorts and a set $\mathcal{F}$ of function symbols, where each $f\in\mathcal{F}$ is attached with its sort declaration $f\colon{\tau_{1}}\times\cdots\times{\tau_{n}}\to\tau_{0}$ ($\tau_{0},\ldots,\tau_{n}\in\mathcal{S}$). The set of sorted variables is denoted by $\mathcal{V}$. For $X\subseteq\mathcal{V}$ and $F\subseteq\mathcal{F}$, the set of terms over the function symbols in $F$ and variables in $X$ is denoted by $\mathcal{T}(F,X)$. We write $t^{\tau}$ to denote that a term $t$ has sort $\tau$. The set of variables appearing in a term $t$ is denoted by $\mathcal{V}\mathsf{ar}(t)$. A term $t$ is said to be linear when any variable occurs at most once in $t$. We denote a sequence of variables by $\vec{x}$, and use the same notation for sequences of function symbols, terms, or similar objects. The set of variables occurring in $\vec{x}$ is denoted by $\{\hskip 0.85358pt\vec{x}\hskip 0.85358pt\}$. Similarly, the set of sequences of elements of a set $T$ is denoted by $T^{*}$.
@@
The set of positions in a term $t$ is denoted by $\mathcal{P}\mathsf{os}(t)$. The symbol and subterm occurring at a position $p\in\mathcal{P}\mathsf{os}(t)$ is denoted by $t(p)$ and $t|_{p}$, respectively. A term obtained from $t$ by replacing subterms at positions ${p_{1}},\dots,{p_{n}}$ by the terms ${t_{1}},\dots,{t_{n}}$, having the same sort as $t|_{p_{1}},\ldots,t|_{p_{n}}$, is written as $t[{t_{1}},\dots,{t_{n}}]_{{p_{1}},\dots,{p_{n}}}$ or just $t[{t_{1}},\dots,{t_{n}}]$ when no confusions arises. If we consider an expression obtained by replacing those subterms $t|_{p_{1}},\ldots,t|_{p_{n}}$ in $t$ by holes of the same sorts, it is called a multihole context and denoted by $t[\penalty 10000\ ]_{{p_{1}},\dots,{p_{n}}}$. A sort-preserving function $\sigma$ from $\mathcal{V}$ to $\mathcal{T}(\mathcal{F},\mathcal{V})$ with a finite domain $\mathcal{D}\mathsf{om}(\sigma)=\{\hskip 0.85358ptx\in\mathcal{V}\mid\sigma(x)\neq x\hskip 0.85358pt\}$ is called a substitution. It is naturally extended to a homomorphism $\sigma\colon\mathcal{T}(\mathcal{F},\mathcal{V})\to\mathcal{T}(\mathcal{F},\mathcal{V})$, which we also denote by $\sigma$. For a set of terms $U$, we write $\sigma(U)=\{\hskip 0.85358pt\sigma(t)\mid t\in U\hskip 0.85358pt\}$. As usual, $\sigma(t)$ can also be also written as $t\sigma$. We write $\{\hskip 0.85358ptx_{1}\mapsto t_{1},\ldots,x_{n}\mapsto\hskip 0.85358pt\}t_{n}$ for the substitution $\sigma$ such that $\sigma(x_{i})=t_{i}$ ($1\leqslant i\leqslant n)$ and $\mathcal{D}\mathsf{om}(\sigma)\subseteq\{\hskip 0.85358ptx_{1},\ldots,x_{n}\hskip 0.85358pt\}$, which may be abbreviated as $\{\hskip 0.85358pt\vec{x}\mapsto\vec{t}\hskip 0.85358pt\}$.
@@
We assume that the set $\mathcal{S}$ of sorts and the set $\mathcal{F}$ of function symbols are partitioned into two disjoint sets, respectively: $\mathcal{S}=\mathcal{S}_{\mathsf{th}}\uplus\mathcal{S}_{\mathsf{te}}$ and $\mathcal{F}=\mathcal{F}_{\mathsf{th}}\uplus\mathcal{F}_{\mathsf{te}}$, where each $f\colon{\tau_{1}}\times\cdots\times{\tau_{n}}\to\tau_{0}\in\mathcal{F}_{\mathsf{th}}$ satisfies $\tau_{i}\in\mathcal{S}_{\mathsf{th}}$ for all $0\leqslant i\leqslant n$. Elements of $\mathcal{S}_{\mathsf{th}}$ ($\mathcal{F}_{\mathsf{th}}$) and $\mathcal{S}_{\mathsf{te}}$ ($\mathcal{F}_{\mathsf{te}}$) are called theory sorts (symbols) and term sorts (symbols). Moreover, we consider a model $\mathcal{M}=\langle\mathcal{I},\mathcal{J}\rangle$ over the sorted signature $\langle\mathcal{S}_{\mathsf{th}},\mathcal{F}_{\mathsf{th}}\rangle$. Here, $\mathcal{I}$ is the interpretation of theory sorts and assigns any sort $\tau\in\mathcal{S}_{\mathsf{th}}$ a non-empty set $\mathcal{I}(\tau)$. Theory symbols are interpreted by $\mathcal{J}$, which assigns any theory symbol $f\colon{\tau_{1}}\times\cdots\times{\tau_{n}}\to\tau_{0}\in\mathcal{F}_{\mathsf{th}}$ to a function $\mathcal{J}(f)\colon\mathcal{I}(\tau_{1})\times\cdots\times\mathcal{I}(\tau_{n})\to\mathcal{I}(\tau_{0})$. A valuation $\rho$ on the model $\mathcal{M}=\langle\mathcal{I},\mathcal{J}\rangle$ is a mapping that assigns any $x^{\tau}\in\mathcal{V}$ with $\rho(x)\in\mathcal{I}(\tau)$. The interpretation of a term $t$ in the model $\mathcal{M}$ over the valuation $\rho$ is denoted by ${t}]\!]_{\mathcal{M},\rho}$.
@@
We further assume that any element $a\in\mathcal{I}(\tau)$ appears as a constant $a^{\tau}$ in $\mathcal{F}_{\mathsf{th}}$, for every sort $\tau\in\mathcal{S}_{\mathsf{th}}$. These constants are called values and the set of all values is denoted by $\mathcal{V}\mathsf{al}$. The set of values contained in a term $t$ is denoted by $\mathcal{V}\mathsf{al}(t)$. A term $t$ is value-free if $\mathcal{V}\mathsf{al}(t)=\varnothing$. For a substitution $\gamma$ we define $\mathcal{V}\mathcal{D}\mathsf{om}(\gamma)=\{\hskip 0.85358ptx\in\mathcal{V}\mid\gamma(x)\in\mathcal{V}\mathsf{al}\hskip 0.85358pt\}$.